Find the equation of the circle which passes through the points (2, –2), and (3, 4) and whose centre lies on the line x + y = 2.

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student
let (h,k) be the centre
as centre lies on the line x+y=2, we get h+k=2........(1)
also equation of circle is (x-h)2+(y-k)2=r2.................(2)
by substituting points (2,-2)and (3,4) in above equation we get 2 different equations
and by solving them with 1 we get the centre (h,k)=(0.7,1.3)
by substituting (h,k) in eqn (2) we get the radius of the circle r=3.55
Hence the equation of circle is(x-0.7)2+(y-1.3)2=3.552
Thanks and Regards
Shaik Aasif
askIITians faculty